# P-NAL: an Effective and Interpretable Entity Alignment Method

### v1

@@ 2.3. A Brief Introduction to NAL
Deduction $A\rightarrow B\ \left\langle f_{1},c_{1}\right\rangle,\ B\rightarrow C\ \left\langle f_{2},c_{2}\right\rangle$ $A\rightarrow C\ \left\langle f=and(f_{1},f_{2}),c=and(f_{1},f_{2},c_{1},c_{2})\right\rangle$
Analogy $A\rightarrow B\ \left\langle f_{1},c_{1}\right\rangle,\ A\leftrightarrow C\ \left\langle f_{2},c_{2}\right\rangle$ $C\rightarrow B\ \left\langle f=and(f_{1},f_{2}),c=and(f_{2},c_{1},c_{2})\right\rangle$
Conditional Deduction $(P\wedge\ Q)\ \Rightarrow R\ \left\langle f_{1},c_{1}\right\rangle,\ Q\ \left\langle f_{2},c_{2}\right\rangle$ $P\Rightarrow R\ \left\langle f=and(f_{1},f_{2}),c=and(f_{1},f_{2},c_{1},c_{2})\right\rangle$
Induction $A\rightarrow B\ \left\langle f_{1},c_{1}\right\rangle,\ A\rightarrow C\ \left\langle f_{2},c_{2}\right\rangle$ $C\rightarrow B\ \left\langle w^{+}=and(f_{2},c_{2},f_{1},c_{1}),w^{-}=and(f_{2},c_{2},not(f_{1}),c_{1})\right\rangle$
Revision $P\ \left\langle f_{1},c_{1}\right\rangle,\ P\ \left\langle f_{2},c_{2}\right\rangle$ $P\ \left\langle w^{+}=w^{+}_{1}+w^{+}_{2},w=w_{1}+w_{2}\right\rangle$
Probabilistic Revision $P\ \left\langle f_{1},c_{1}\right\rangle,\ P\ \left\langle f_{2},c_{2}\right\rangle$ $P\ \left\langle f=or(f_{1},f_{2}),w=w_{1}+w_{2}\right\rangle$
2.4. Why NAL
Actually there might be many different logical systems that are qualified to represent the similarity inference process of EA. However, we believe that the non-axiomatic nature of NAL fits in the domain of knowledge graph better than those axiomatic logical systems, because real world KGs need to deal with the problem of open-domain and alterable/incomplete/conflicting facts. Fundamentally, the tasks of knowledge graph (such as EA), fits well with the assumption of insufficient knowledge and resources  (Wang, 2013), which is the basic assumption of NAL.
3. Our Approach
Our EA approach’s main idea stems from the well-known EA method PARIS. The overall structure of P-NAL adopts an iterative aligning strategy, and for each iteration it first performs similarity inference, then it uses a matching technique (rBMat algorithm with modification in Section 3.4) to obtain EA results. We formalize the similarity inference as using NAL’s revision inference rule to aggregate two types of inference paths. P-NAL also infers the matching of relations in each iteration. We explore and implement some other design choices and tricks to complete the alignment framework, as illustrated in the following subsections.
Figure 1. An illustration of align-bridge, omitting irrelevant triples.
3.1. PARIS Reinterpreted in the Formal Language of NAL
The key point of the alignment process in PARIS is finding instances of bridge-like inference path between to-be-aligned entity pairs. We coin the term “align-bridge” to refer to such type of path. Valid align-bridges are retrieved from the KGs in a depth-first manner. As shown in Figure 1, entity $y_{1}$ and entity $y_{2}$ belongs to different KGs (where subscripts represent different KGs) and ($y_{1}$, $y_{2}$) forms a to-be-aligned entity pair. Triple ($x_{1}$$r_{1}$$y_{1}$) is a relational or attribute triple in $\mathcal{KG}_{1}$, where $x_{1}$ is either an entity or a literal respectively. Note that PARIS (and P-NAL) automatically duplicates every original KG triple ($a$$b$$c$) with a reversed triple ($c$$b^{-1}$$a$) upon KG loading, so the attribute triple ($x_{1}$$r_{1}$$y_{1}$) with a literal $x_{1}$ is a reversed attribute triple. Similarly, triple ($x_{2}$$r_{2}$$y_{2}$) is a relational or attribute triple in $\mathcal{KG}_{2}$.
In this paper, every entity, literal or relation is regarded as an atomic term in NAL. Triple ($x_{1}$$r_{1}$$y_{1}$) is reinterpreted as inheritance statement (*, $x_{1}$$y_{1}$) $\rightarrow$ $r_{1}$. Its intuitive meaning is “The relation between $x_{1}$ and $y_{1}$ is a specialization of relational term $r_{1}$”. The triples (or “facts”) of the KGs can be seen as absolutely true to some extent, so the truth-value attached to the statement is $\left\langle 1,1\right\rangle$. All of the triples of the two KGs are taken as input sentences, which forms the experience of the inference system. In PARIS, the equality score between $x_{1}$ and $x_{2}$ is retrieved to measure the similarity between them. We interpret the equality as a similarity statement $x_{1}$leftrightarrow  $x_{2}$ and the score is reflected in the truth-value of the statement. Note that in the case of entity pair, the similarity comes from either seed alignments or alignments of the previous iteration. We omit the entity similarity statement which has a $f$ or $c$ lesser than $theta$, a hyper-parameter. And in the case of literal pair, the similarity comes from literal comparison (for example, string identity comparison).
At the same time, PARIS retrieves the sub-relation probability score Pr($r_{1}$subseteq $r_{2}$) (from the computation result of the previous iteration or a default value $iota$, which is a hyper-parameter) which we interpret as an inheritance statement $r_{1}$rightarrow $r_{2}$ with a truth-value. Its intuitive meaning is “The relational term $r_{1}$ is a specialization of relational term $r_{2}$”. PARIS evaluates the degree of functionality of relation $r_{2}$ with precomputed functionalities of each relation. We interpret it as an inheritance statement $r_{2}$rightarrow [$fun$] with the degree reflected in the truth-value. The statement intuitively means “$r_{2}$ has the functional property (to some extent)”.
Then the validity of align-bridge is interpreted in terms of inference steps in NAL as follows:
$(type\ I\ path)\ premises:$
(1) $\left(*,x_{1},y_{1}\right)\rightarrow r_{1}\ \ \left\langle 1,1\right\rangle.$
(2) $\left(*,x_{2},y_{2}\right)\rightarrow r_{2}\ \ \left\langle 1,1\right\rangle.$
(3) $r_{1}\rightarrow r_{2}\ \ \left\langle f_{3},c_{3}\right\rangle.$
(4) $r_{2}\rightarrow r_{1}\ \ \left\langle f_{4},c_{4}\right\rangle.$
(5) $x_{1}\leftrightarrow x_{2}\ \ \left\langle f_{5},c_{5}\right\rangle.$
(6) $r_{1}\rightarrow[fun]\ \ \left\langle f_{6},c_{6}\right\rangle.$
(7) $r_{2}\rightarrow[fun]\ \ \left\langle f_{7},c_{7}\right\rangle.$
((,\$$(\left(*,\$a,\#b_{1}\right)\rightarrow\$r\ \wedge\ \left(*,\$a,\#b_{2}\right)\rightarrow\$r$
(8) \$$\wedge\ \$r\rightarrow[fun])\ \Rightarrow\#b_{1}\leftrightarrow\#b_{2}\ \ \left\langle f_{8},c_{8}\right\rangle.$
$inference\ steps\ (path)\ \&\ conclusion:$
(9) $from\ (1)\ and\ (3),Deduction:\left(*,x_{1},y_{1}\right)\rightarrow r_{2}\ \ \left\langle f_{9},c_{9}\right\rangle.$
(10) $from\ (2)\ and\ (5),Analogy:\left(*,x_{1},y_{2}\right)\rightarrow r_{2}\ \ \left\langle f_{10},c_{10}\right\rangle.$
(11) $from(8),(9),(10),and(7),Conditional\ deduction\times 3:\ \ y_{1}\leftrightarrow y_{2}\ \ \left\langle f_{11},c_{11}\right\rangle.$
Here is a brief introduction of the idea of the inference steps: The first two steps aim to match up the two triples (1) and (2). The first step exchanges $r_{1}$ in the triple (1) for $r_{2}$. The second step exchanges $x_{2}$ in the triple (2) for $x_{1}$. The third step is a type of inference similar with the resolution rule of Horn clauses and (8) correspond to a Horn rule.
@@ 3.1. PARIS Reinterpreted in the Formal Language of NAL
In the path listed above, we omit two auxiliary inference steps right before arriving at conclusion (10) which performs structural transformation in order to dismount $x_{2}$ from the product without modifying truth-value. The last conditional deduction of (11) degenerates into a case without conjunction in its premises (similar with Modus Ponens) and its truth function remains the same. Statement (11) is the conclusion of the above inference steps and the whole steps act as a summarizing or validation process of the align-bridge. Note that there can be four distinct inference paths of an align-bridge (including the path above) that share the same premise set and the same conclusion (with slightly different truth-values), with different details of inference steps. In implementation, we aggregate two of them (of both relational inheritance direction) by probabilistic revision rule (similar with PARIS).
Implication statement (8) is regarded as a definition or a piece of essence of the concept “functionality”. Relations’ functionality seems to reflect a widespread orderliness of reality or human cognition and PARIS leverages such orderliness.
The conclusions with the same statement but obtained from different align-bridges are merged by probabilistic revision rule because of the probabilistic nature of functionality. For example, the functionality of relation $children$ is 0.68 which means that the majority of the population approximately have one to three children. Entity $Lynne\_Che\-ney$ has two children, however, in the alignment system when performing alignment of the two $Mary\_Cheney$ entities (who is a children of $Lynne\_\-Cheney$) of different KGs we do not know how many children does $Lynne\_Cheney$ has. The conclusion of the align-bridge $Mary\_\-Chen\-ey(en)$$Lynne\_Cheney(en)$$Lynne\_Cheney(zh)$$Ma\-ry\_Cheney\-(zh)$ is
$Mary\_Cheney(en)\leftrightarrow Mary\_Cheney(zh)\ \ \left\langle f,c\right\rangle.$
It has a probabilistic nature because we don’t know whether $Mary\_\-Cheney(en)$ and $Mary\_Cheney(zh)$ is the same children of $Lynne\_\-Cheney$. Thus the probabilistic revision rule is used to aggregate the conclusions of multiple align-bridges. This rule is similar with the continued multiplication of PARIS’s probability formula for $Pr\left(x\equiv x^{\prime}\right)$ (given in Section 2.2), except for the introduction and calculation of confidence.
3.2. Aligning Relations
Apart from aligning entities, we also align relations using an adapted version of PARIS’s probabilistic relation aligning method. The relation aligning formula of PARIS is reinterpreted as directly estimating the evidence amount for the conclusion. The truth-value of the conclusion statement $r_{1}\rightarrow r_{2}$ (the inheritance statement between two relations $r_{1}$ and $r_{2}$) is computed as follow:
$w^{+}=\Sigma_{(x_{1},r_{1},y_{1})}\left(1-\prod_{(x_{2},r_{2},y_{2})}\left(1-expt\left(x_{1}\leftrightarrow x_{2}\right)\times expt\left(y_{1}\leftrightarrow y_{2}\right)\right)\right)$
$w=\Sigma_{(x_{1},r_{1},y_{1})}\left(1-\prod_{x_{2},y_{2}}\left(1-expt\left(x_{1}\leftrightarrow x_{2}\right)\times expt\left(y_{1}\leftrightarrow y_{2}\right)\right)\right)$
where $expt$ is the $expectation$ of a truth-value, and $expectation$ is a combined measurement of $f$ and $c$, defined as $expectation=f\times c$.
Alternatively, the truth-value of statement $r_{1}\rightarrow r_{2}$ can be obtained via inference on two types of paths:
$(type\ III\ path)\ premises:$
(12) $\left(*,x_{1},y_{1}\right)\rightarrow r_{1}\ \ \left\langle 1,1\right\rangle.$
(13) $\left(*,x_{2},y_{2}\right)\rightarrow r_{2}\ \ \left\langle 1,1\right\rangle.$
(14) $x_{1}\leftrightarrow x_{2}\ \ \left\langle f_{14},c_{14}\right\rangle.$
(15) $y_{1}\leftrightarrow y_{2}\ \ \left\langle f_{15},c_{15}\right\rangle.$
$inference\ steps\ (path)\ \&\ conclusion:$
(16) $from\ (12)\ and\ (14),Analogy:\left(*,x_{2},y_{1}\right)\rightarrow r_{1}\ \ \left\langle f_{16},c_{16}\right\rangle.$
(17) $from\ (16)\ and\ (15),Analogy:\left(*,x_{2},y_{2}\right)\rightarrow r_{1}\ \ \left\langle f_{17},c_{17}\right\rangle.$
(18) $from\ (13)\ and\ (17),Induction:r_{1}\rightarrow r_{2}\ \ \left\langle 1,c_{18}\right\rangle.$
$(type\ IV\ path)\ premises:$
(19) $\left(*,x_{1},y_{1}\right)\rightarrow r_{1}\ \ \left\langle 1,1\right\rangle.$
(20) $\left(*,x_{2},y_{2}\right)\rightarrow r_{2}\ \ \left\langle 0,1\right\rangle.$
(21) $x_{1}\leftrightarrow x_{2}\ \ \left\langle f_{21},c_{21}\right\rangle.$
(22) $y_{1}\leftrightarrow y_{2}\ \ \left\langle f_{22},c_{22}\right\rangle.$
$inference\ steps\ (path)\ \&\ conclusion:$
(23) $from\ (19)\ and\ (21),Analogy:\left(*,x_{2},y_{1}\right)\rightarrow r_{1}\ \ \left\langle f_{23},c_{23}\right\rangle.$
(24) $from\ (23)\ and\ (22),Analogy:\left(*,x_{2},y_{2}\right)\rightarrow r_{1}\ \ \left\langle f_{24},c_{24}\right\rangle.$
(25) $from\ (20)\ and\ (24),Induction:r_{1}\rightarrow r_{2}\ \ \left\langle 0,c_{25}\right\rangle.$
Premise (20) has a frequency value of 0, so it is not a triple in the KG but represents the absence of the triple. Note that the induction inference rule is a weak inference rule, so the upper bound of its conclusion’s confidence is lower than the strong inference rules (such as deduction and analogy). Type III path only generates positive evidence for the conclusion and type IV path only generates negative evidence, because of the characteristic of the truth-function of induction rule. The conclusions of the two types of path are supposed to be merged by the revision rule. In implementation, for simplicity we directly estimate evidence amount instead of performing inference for relation alignment, however, the latter approach may provide a more theoretical view. These two approaches share the same idea but the calculation processes are not identical.
Due to the incompleteness of the KGs of relevant datasets, the positive evidence for the inheritance statement $r_{1}\rightarrow r_{2}$ is usually inadequate, for that either of the premise (12) or (13) may be missing in the KG but actually true. To address this issue, we force the frequency of every inheritance statement to increase by a proportion (at the end of each iteration), that is $f:=f+\left(1-f\right)\times inc\_relation\_f$, where :=assign:=:= represents assignment and $inc\_relation\_f$ is a hyper-parameter.
3.3. Utilizing Literal Values
@@
Literal values in real-world KGs act as entity names, entity descriptions, relation/attribute names or attribute values, carrying enormous information. Literal values include texts (strings), numerical values and dates. Literals’ deficiency of its outer semantic structure (triples) contrasts with its abundant internal semantics. However, symbolic reasoning languages (systems) like NAL currently can’t effectively handle the subtle semantics in texts for the following reasons: semantic parsing/understanding requires processing capacity and efficiency of complex logical forms and it also requires automatic learning capacity; the lacking of KGs with complex logical forms; the lacking of KGs with detailed and comprehensive common sense knowledge. In a certain perspective, the literal values in real-world KGs are not really “literal” but rather under-characterized entities, concepts, triples, common sense knowledge and/or statements with complex logical forms. The real-world KG project may not have enough information or adequate paradigm to deal with them. For example, the literal value of attribute triple ((((John Lennon, deathPlace, “Manhattan, New York City, United States”@en)))) referred to entities “Manhattan”, “New York” and “United States”, and its form indicates a specific relation between these places.
Deep neural network language models provide an interim solution to the literal value understanding problem. For example, BERT-INT  (Tang et al., 2020) utilize BERT to embed names/descriptions and values into vector space, thus use similarities between the feature vectors for alignment.
P-NAL adopts the same embedding method as BERT-INT and fuse it into the similarity inference system. First, the basic BERT unit is finetuned on the name/description of seed alignment entity pairs. Then we use the finetuned model to compute feature vectors for all entities (name/description). Moreover, we get the feature vectors for attribute values. Entity name/description feature’s cosine similarity is used to convert directly to the statement
$y_{1}\leftrightarrow y_{2}\ \ \left\langle sim(y_{1},y_{2}),\ C_{name}\right\rangle.$
where $sim$ is cosine similarity and $C_{name}$ is a hyper-parameter. The direct path linking the to-be-aligned entities with their name/description similarity is the type II path of similarity inference as mentioned before. The statement is seen as a piece of evidence and fused with other evidences (such as those from align-bridges) by revision rule.

### v2

@@ Appendix A Introduction of NAL
Deduction $A\rightarrow B\ \left\langle f_{1},c_{1}\right\rangle$ $A\rightarrow C\ \left\langle f=and(f_{1},f_{2}),c=and(f_{1},f_{2},c_{1},c_{2})\right\rangle$
$B\rightarrow C\ \left\langle f_{2},c_{2}\right\rangle$
Analogy $A\rightarrow B\ \left\langle f_{1},c_{1}\right\rangle$ $C\rightarrow B\ \left\langle f=and(f_{1},f_{2}),c=and(f_{2},c_{1},c_{2})\right\rangle$
$A\leftrightarrow C\ \left\langle f_{2},c_{2}\right\rangle$
Conditional $(P\wedge Q)\Rightarrow R\left\langle f_{1},c_{1}\right\rangle$ $P\Rightarrow R\ \left\langle f=and(f_{1},f_{2}),c=and(f_{1},f_{2},c_{1},c_{2})\right\rangle$
Deduction $Q\ \left\langle f_{2},c_{2}\right\rangle$
Induction $A\rightarrow B\ \left\langle f_{1},c_{1}\right\rangle$ $C\rightarrow B\left\langle w^{+}=and(f_{2},c_{2},f_{1},c_{1}),w^{-}=and(f_{2},c_{2},not(f_{1}),c_{1})\right\rangle$
$A\rightarrow C\ \left\langle f_{2},c_{2}\right\rangle$
Revision $P\ \left\langle f_{1},c_{1}\right\rangle$ $P\ \left\langle w^{+}=w^{+}_{1}+w^{+}_{2},w=w_{1}+w_{2}\right\rangle$
$P\ \left\langle f_{2},c_{2}\right\rangle$
Probabilistic $P\ \left\langle f_{1},c_{1}\right\rangle$ $P\ \left\langle f=or(f_{1},f_{2}),w=w_{1}+w_{2}\right\rangle$
Revision $P\ \left\langle f_{2},c_{2}\right\rangle$
Table 2: The table of relevant rules and their truth functions.
Appendix B Related Work
B.1 Embedding-based EA
Embedding-based EA methods usually consist of three parts: the embedding module, the alignment module and the matching module. For the embedding module, translational methods and graph neural network (GNN) methods are the most popular. Translational methods, such as MTransE  Chen et al. (2016), usually optimize a margin-based loss function to learn the structural information (relation triples) of a KG. On the other hand, GNN methods recursively aggregate the representations of neighboring nodes with graph convolutional networks (GCNs) or graph attention networks (GATs). The representative ones are RDGCN  Wu et al. (2019) and RREA  Mao et al. (2020b), respectively. The alignment module maps the entity embeddings in different KGs into a unified space. There are generally three techniques  Fanourakis et al. (2023) for this module: 1. Sharing the embedding space by using the margin-based loss to enforce the seed alignment entities’ embeddings from different KGs to be close. 2. Swapping the triples of seed alignment entities. 3. Mapping the entity vectors from one embedding space to the other using a transformation matrix. The matching module generates the final alignment result. Common practices use the cosine similarity, the Manhattan distance, or the Euclidean distance between entity embeddings to measure their similarities and then perform a specific matching algorithm based on the similarity scores.
B.2 Path-based EA
PARIS  Suchanek et al. (2011) is a classic unsupervised non-neural EA method with competitive performance on benchmark datasets  Leone et al. (2022). It is purely path-based. Following previous works  Hogan et al. (2010), PARIS introduces the probabilistic usage of "functionality" into the field of EA to enhance the validity of similarity inference paths. Functionality generally corresponds to the uniqueness of related things, for example a man can only have one father but multiple friends, so $fun(father)$ is close to 1 and $fun(friend)$ is relatively lower, where $fun()$ represents functionality of a relation or attribute. See  Suchanek et al. (2011) for more details about functionality. With functionality, PARIS constructs a probabilistic model that estimates the probabilities of entity equivalences:
$Pr\left(x_{1}\equiv x_{2}\right)=1-\prod_{r_{1}(x_{1},y_{1}),r_{2}(x_{2},y_{2})}(1-Pr\left(r_{1}\subset r_{2}\right)\times fun(r_{2}^{-1})\times$ $Pr\left(y_{1}\equiv y_{2}\right))$
As depicted in the above formula, PARIS estimates the equivalence probabilities by integrating paths that connects corresponding entities. It also find subrelations between the two ontologies of KG. Subrelations, such as $r_{1}\subset r_{2}$, intuitively means a correspondence of two relations of different KGs such that one relational fact of $r_{1}$ in $\mathcal{KG}_{1}$ implies the existence of a corresponding relational fact of $r_{2}$ in $\mathcal{KG}_{2}$. Here is the formula for $Pr\left(r_{1}\subset r_{2}\right)$:
$\frac{\Sigma_{r_{1}(x_{1},y_{1})}\left(1-\prod_{r_{2}(x_{2},y_{2})}\left(1-Pr\left(x_{1}\equiv x_{2}\right)\times Pr\left(y_{1}\equiv y_{2}\right)\right)\right)}{\Sigma_{r_{1}(x_{1},y_{1})}\left(1-\prod_{x_{2},y_{2}}\left(1-Pr\left(x_{1}\equiv x_{2}\right)\times Pr\left(y_{1}\equiv y_{2}\right)\right)\right)}$
With the help of subrelations’ measurement, PARIS generalizes the equation of $Pr\left(x_{1}\equiv x_{2}\right)$ to the case where the two ontologies do not share common relations. Therefore, PARIS recursively aligns the entities and the equivalence probability of $x_{1}\equiv x_{2}$ depends recursively on other equivalence probabilities. In each iteration, the probabilities are re-calculated based on the equivalences and subrelations of the previous iteration. Initial equivalences are computed between attribute literals based on a certain string distance measurement.
PARIS+  Leone et al. (2022) is a variant of PARIS that makes a simple refinement and works in the absence of attribute triples. It processes the seed alignment information to generate synthetic attribute triples. That is, for every pair of seed alignments ($x_{1}$, $x_{2}$), it creates the attribute triples ($x_{1}$, EA:label, string($x_{1}$)) and ($x_{2}$, EA:label, string($x_{1}$)), where EA:label is a synthetic relation. Thus, the reverse of the relation EA:label is designed to be highly functional in order to let the model match the seed alignments easily. NALA adopts the same refinement as PARIS+.
B.3 Combined EA
BERT-INT  Tang et al. (2020), an embedding-path EA method, uses the well-known transformer model BERT to embed the entities and literals. It calculates the cosine similarity of the entity name/description embedding. Then it proposes an interaction model that compares each pair of neighbors or attributes (which forms a path from the source entity to the target entity) to obtain the neighbor/attribute similarity score. The name/description similarity vector, neighbor similarity vector and attribute similarity vector are concatenated and applied to a MLP layer to get the final similarity score.
FGWEA  Tang et al. (2023) is a three-step progressive optimization algorithm for EA and it can be classified as an embedding-path EA method. First, the entity names and concatenated attribute triples are used for semantic embedding matching to obtain initial anchors. Then in order to approximate GWD (Gromov-Wasserstein Distance  Peyré et al. (2016)), FGWEA computes cross-KG structural and relational similarities, which are then used for iterative multi-view optimal transport alignment. Finally, the Bregman Proximal Gradient algorithm  Xu et al. (2019) is employed to refine the GWD’s coupling matrix.
B.4 Other EA Methods
There are also a few works that focus on the interpretability or explanability of EA, such as LightEA  Mao et al. (2022) and ExEA  Tian et al. . LightEA is an interpretable non-neural EA method. It is inspired by a classical graph algorithm, label propagation  Zhu and Zoubin (2002). First, it generates a random orthogonal label for each seed alignment entity pair. Then, the labels of entities and relations are propagated according to the three views of adjacency tensor. Finally, LightEA utilizes sparse sinkhorn iteration to address the assignment problem of alignment results.
The ExEA framework, proposed by  Tian et al. , aims to explain the results of embedding-based EA. It generates semantic matching subgraphs as explanation by matching semantically consistent triples around the two aligned entities. ExEA devises an alignment dependency graph structure to gain deeper insights into the explanation.
The recent literature of EA is abundant, focusing on many different aspects or procedures of entity alignment apart from the aforementioned ones, such as utilizing attribute triples  Liu et al. (2020); Sun et al. (2017), utilizing literals  Gesese et al. (2021); Chen et al. (2018) , sample mining  Liu et al. (2022); Mao et al. (2021a), reinforcement learning  Guo et al. (2022), matching algorithm  Lin et al. (2023); Dao et al. (2023); Mao et al. (2021b); Xu et al. (2020); Zeng et al. (2020), iterative strategy  Liu et al. (2023); Mao et al. (2020a) and unsupervised learning  Jiang et al. (2023a, b); Liu et al. (2022); Luo and Yu (2022); Zhao et al. (2022). There are also some surveys for EA  Fanourakis et al. (2023); Zeng et al. (2021); Sun et al. (2020); Mao et al. (2022). Besides graph structural, attribute and literal information, there are other information forms researched by the EA community, such as temporal, spatial and graphical information, however, these topics are beyond the scope of this paper.
Appendix C Additional Remarks on Inference Paths
@@
As for type I path, we omit two auxiliary inference steps right before arriving at conclusion (6) which performs structural transformation in order to dismount $x_{2}$ from the product of (4) without modifying its truth-value. The conditional deduction of (11) degenerates into a case without conjunction in its premises (similar with Modus Ponens) and its truth function remains the same. Note that in the path only one direction of the relational inheritance is considered ($r_{1}\rightarrow r_{2}$) and there exists a symmetrical variation of the path that utilizes the other direction ($r_{2}\rightarrow r_{1}$). The conclusions of the two paths are aggregated by probabilistic revision rule.
Appendix D Discussion
D.1 Problem of Understanding Literal Value
@@
Literal values in real-world KGs act as entity names, entity descriptions, relation/attribute names or attribute values, carrying enormous information. Literal values include texts (strings), numerical values and dates. Deep neural network language models provide an interim solution to the problem of understanding literal values. For example, BERT-INT  Tang et al. (2020) utilize BERT to embed names/descriptions and values into vector space, thus use similarities between the feature vectors for alignment. Literals’ deficiency of its outer semantic structure (triples) contrasts with its abundant internal semantics. However, symbolic reasoning languages (systems) like NAL currently can’t effectively handle the subtle semantics in texts for the following reasons: semantic parsing or understanding requires processing capacity and efficiency of complex logical forms and it also requires automatic learning capacity; the lacking of KGs with complex logical forms; the lacking of KGs with detailed and comprehensive common sense knowledge. In a certain perspective, the literal values in real-world KGs are not really "literal" but rather under-characterized entities, concepts, triples, common sense knowledge and/or statements with complex logical forms. The real-world KG project may not have enough information or adequate paradigm to deal with them. For example, the literal value of attribute triple ((((John Lennon, deathPlace, "Manhattan, New York City, United States"@en)))) referred to entities "Manhattan", "New York" and "United States", and its form indicates a specific relation between these places.
D.2 Understanding of type II path